OpenStudy (anonymous):

I have to prove that sqrt 11 is irrational, which I have done, ..Is sqrt11 + 3 irrational, please prove?????? Does my head in I tell you :)

OpenStudy (zarkon):

\[\sqrt{11}+3\]?

OpenStudy (zarkon):

prove it is irrational by using proof by contradiction

OpenStudy (anonymous):

I am a bit lost with the plus 3 new to this

OpenStudy (zarkon):

suppose \(\sqrt{11}+3\) is rational

OpenStudy (zarkon):

then \[\sqrt{11}+3=\frac{a}{b}\] a,b integers

OpenStudy (zarkon):

then \[\sqrt{11}=\frac{a}{b}-3\]

OpenStudy (anonymous):

still with you :)

OpenStudy (zarkon):

\[\frac{a}{b}-3\] is a rational number but \[\sqrt{11}\] is not this is a contradiction thus \(\sqrt{11}+3\) is irrational

OpenStudy (anonymous):

Is that all I need to say?

OpenStudy (zarkon):

what else would you need to add?

OpenStudy (anonymous):

ahhhhh because regardless you are still dealing with an irrational number and nothing will change that?

OpenStudy (anonymous):

unless you had alot of time on your hands LOL

OpenStudy (anonymous):

Is this a correct assumption?